OpenStudy (anonymous):

find the exact value of cos^2(pi/8) using appropriate power reducing formula! WILL MEDAL

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\frac{\pi}{8}=\frac{\pi}{4}\times \frac{1}{2}\]so you can use the 'half angle" formula \[\cos(\frac{\pi}{8})=\sqrt{\frac{1+\cos(\frac{\pi}{4})}{2}}\]

OpenStudy (misty1212):

even easier for you, since you want \[\cos^2(\frac{\pi}{8})\] just ignore the square root symbol

OpenStudy (misty1212):

thanks for the medal @freckles \(\color\magenta\heartsuit\) \(\color\magenta\heartsuit\) \(\color\magenta\heartsuit\)

OpenStudy (misty1212):

@wio too!! \(\color\magenta\heartsuit\) \(\color\magenta\heartsuit\)